1). The average (arithmetic mean) of 7 different numbers is 5

Consider the following two quantities:

A. The median of the 7 numbers. B. 5.

A. Quantity A is greater than Quantity B

B. Quantity B is greater than Quantity A

C. The two quantities are equal.

D. The relationship cannot be determined without more information.

2). List S = {16, 9, 23, X, 13, 16}

In list S, the mean and median are equal to one another. What is the value of X?
Answer #1

1)

D. The relationship cannot be determined without more information.

2)

for sum of all values =77+x

as 16 falls apprximately in midddle if we put all values in ascending order ; therefore it should be median

hence x+77=6*16

x =96-77 =19